I am planning to purchase tickets for large group of people.
I have to end up with 105 tickets total. For every 15 tickets we buy, we get one free.
What is the formula or equation to calculate this type of "buy 15, get 1 free" problem.
Thanks, G. Brown

Answer by checkley77(12844)

105/(15+1)
105/16=6.5625
THUS YOU NEED TO BY 6 "BUY 15 GET 1 FREE" TICKETS.
6*16=96
105-96=9 INDIVIDUAL TICKETS ALSO NEED TO BE PURCHASED VTO MAKE THE 105.